OpenStudy (anonymous):

what is the lcd of 8 and 3

OpenStudy (anonymous):

Unfortunately, it's just 1, because they're not divisible by anything else.

OpenStudy (anonymous):

Wait. I take that back. I was thinking of factors. The correct answer is 24, as that's the smallest number that both are able to divide into.

OpenStudy (anonymous):

24

OpenStudy (anonymous):

well, is it twenty four? ;P

OpenStudy (anonymous):

For these I find it's best to factor the numbers in question: \[8 = 2^3\]\[3=3^1\] So to get a common denominator, we need the union of these factors (and their multiplicities) \[LCD = 2^3\cdot 3 = 8 \cdot 3 = 24\] In this case it's trivial to just multiply them, but it can make things easier if they have some factors in common..